Two moles of a monatomic ideal gas are at a temperature of 309 5. The gas expands reversibly and isothermally to twice its original volume. Calculate the work done by the gas, the heat supplied and the change in the internal energy


1
Expert's answer
2021-03-24T19:40:08-0400

"A=\nu RT \text{ln}\frac{V_2}{V_1}=2\cdot 8.31\cdot 582\cdot \text {ln}2=6.7~\text{kJ},"

"Q=A=6.7~\text{kJ},"

"\Delta U=0~\text{kJ}."